# Text2VDM: Text to Vector Displacement Maps for Expressive and Interactive 3D Sculpting

### v1

@@ 4 Experiments
In this section, we conduct experiments to evaluate the various capabilities of Text2VDM both quantitatively and qualitatively for text-to-VDM brush generation. We then present an ablation study that validates the significance of our key insight into CFG-weighted SDS, as well as the effect of the region control and shape control.
@@ 4.1 Qualitative Evaluation
To the best of our knowledge, Text2VDM is the first framework to generate VDM brushes from text. We adapted three existing methods for comparison and classified them into two categories. The first category includes Text2Mesh [28] and TextDeformer [10], which generate a brush mesh through text-guided mesh deformation on a planar mesh, following a process similar to ours. For the second category, we opt to directly generate VDM via Paint-it [56]. Notably, this method originally uses SDS to optimize a UNet for generating PBR textures. We reframed it to suit our VDM brush generation task, modifying it to generate VDM through SDS optimization of the UNet. We compared the visual results in Figure 5 and Figure 6.
Compared to other methods, Text2VDM can generate more vivid and better-quality VDM brushes. Text2Mesh applies displacement to each vertex along normal directions, resulting in limited mesh deformation. TextDeformer indicates the accumulation of local deformations in the Jacobians, which results in global mesh drift, making it challenging to bake these meshes into VDM. Reframed Paint-it VDM generation is equivalent to optimizing the three-axis displacement of each vertex on the mesh with SDS. Although the UNet reduces noise from the SDS [56], geometric regularization is still required to ensure mesh quality. The generated mesh must compromise between solving the problem and maintaining smoothness, which makes achieving high-quality mesh generation quite challenging.
@@ 4.2 Quantitative Evaluation
Figure 7: Effect of CFG-weighted SDS. CFG-weighted SDS effectively mitigates semantic coupling issues in SDS, such as generating the tortoise’s tail and head or the snail’s head, by providing more focused semantic guidance. In contrast, CSD adds extra negative terms that fail to decouple semantics, resulting in a less stable and more time-consuming optimization process.
@@
User Study. We further conducted a user study to evaluate the effectiveness and expressiveness of our method. A Google Form was utilized to assess 1) geometry quality and 2) consistency with text. We recruited 32 participants, of whom 14 are graduate students majoring in media arts, and 18 are company employees specializing in AI content generation. In this form, the participants were instructed to choose the preferred renderings of VDM from different methods in randomized order, as shown in Table 2. The results show participants preferred our method by a significant margin.
@@ 4.3 Ablation Study
Effects of CFG-Weighted SDS. We conducted experiments to compare the generated results of directly using SDS [34], CFG-weighted SDS, and CSD [57] with three different annealed weights of negative prompt (Figure 7). As discussed in Section 3.3, SDS can result in semantic coupling when generating sub-object structure, leading to artifacts like the tortoise’s tail and head or the snail’s head. We also found that using negative prompts was ineffective at decoupling semantics. Increasing the initial weight of negative prompts further makes the optimization unstable, resulting in low-quality results. In contrast, our method effectively mitigates semantic coupling to produce high-quality meshes without requiring additional UNet inference.
Figure 8: Effect of region control. Region masks can effectively control the shape of surface details based on different text inputs.
Effects of Region Control. Figure 8 demonstrates two sets of region masks and their control over surface details generation under different text prompts. Without using a region mask, the results lack a specific shape, which may not satisfy the desired stylized effect. By using a region mask, our generated results effectively conform to the user’s desired shapes while also aligning with the styles specified by the text, such as metal and stone.
Effects of Shape Control. Our method demonstrates that user-specified VDMs can effectively control the volume and direction of generated geometric structures. As shown in Figure 9, various generated geometric structures, such as elf ears and pauldrons, are high-quality and align with the text descriptions. We also found that without volume initialization, it is challenging to generate desired results. It indicates that this initialization is crucial for steering the gradient flow of geometric structure generation via adjusting the Laplacian term.
Figure 9: Effect of shape control. User-specified VDMs can help achieve the intended final effect of geometric structures by initializing the brush’s volume and direction.
@@ 4.4 Applications
Once various VDM brushes are generated, users can directly use these brushes to meet diverse creative needs in mainstream modeling software. For example, they can apply VDM brushes for mesh stylization and engage in a real-time iterative modeling process.
Local-to-Global Mesh Stylization. Although mesh stylization is a complex task even for professional artists, combining different surface details allows users to achieve stylization quickly. For instance, users can apply a variety of wall-damage brushes to specific areas of a stone pillar, creating a style of damage ( Figure 10).
Figure 10: Local to global mesh stylization. Applying various surface details brushes can create a damaged-style stone pillar model.
Coarse-to-Fine Interactive Modeling. Unlike previous methods [4, 58] that require a lengthy optimization process for each edit and result in non-reusable outcomes, our generated VDM brushes can be directly used in modeling software. This enables users to apply the generated brushes easily and interactively. For example, Figure 11 shows that users can combine various brushes, such as skeleton hand, rose pattern, and pauldron to refine a coarse cloth model into a highly detailed one.
@@ 5 Conclusion
We have presented Text2VDM, a novel framework for VDM brush generation from text. A VDM is a non-natural 2D image where each pixel stores a 3D displacement vector, making it challenging for existing T2I models to generate. Thus, We treat VDM generation as mesh deformation via the Laplace-Beltrami operator from a dense planar mesh. To generate the intended effects of surface details and geometric structures, we provide two control methods: region control and shape control. Additionally, VDM brushes often contain sub-object structures, which can lead to semantic coupling issues in SDS. We propose using CFG-weighted blending for prompt tokens to effectively mitigate this, achieving high-quality brush generation. The generated VDM brushes are directly compatible with mainstream modeling software, enabling various applications such as mesh stylization and real-time interactive modeling.
Limitations and future work. While our framework can generated high-quality VDM brushes, they may encounter multi-view inconsistencies, a common issue introduced by SDS. To further address it, the view-consistent diffusion model proposed in MV2MV [6] may be helpful.

### v2

@@ 4 Experiments
We conducted experiments to evaluate the various capabilities of Text2VDM both quantitatively and qualitatively for text-to-VDM brush generation. We then present an ablation study that validates the significance of our key insight into Semantic Enhancement SDS, as well as the effect of the region mask and VDM initialization.
@@ 4.1 Qualitative Evaluation
To the best of our knowledge, Text2VDM is the first framework to generate VDM brushes from text. We adapted three existing methods for comparison and classified them into two categories. The first category includes Text2Mesh [31] and TextDeformer [11], which generate a brush mesh through text-guided mesh deformation on a planar mesh, following a process similar to ours. For the second category, we opt to directly generate VDM via Paint-it [58]. Notably, this method originally uses SDS to optimize a UNet for generating PBR textures. We reframed it to suit our VDM brush generation task, modifying it to generate VDM through SDS optimization of the UNet. In geometric structures generation experiment (Figure 6), all methods are compared fairly, with the same non-zero VDM initialization and mask applied to each prompt (see Appendix B.2). For surface details (Figure 5), all methods start with a zero-valued VDM and no masks to ensure a fair comparison.
Compared to other methods, Text2VDM can generate better-quality VDM brushes. Text2Mesh applies displacement to each vertex along normal directions, resulting in limited mesh deformation. TextDeformer indicates the accumulation of local deformations in the Jacobians, which results in global mesh drift, making it challenging to bake these meshes into VDM. Reframed Paint-it VDM generation is equivalent to optimizing the three-axis displacement of each vertex on the mesh with SDS. Although the UNet reduces noise from the SDS [58], smooth regularization is still required to ensure mesh quality, which makes achieving high-quality mesh generation quite challenging.
@@ 4.2 Quantitative Evaluation
Figure 7: Effect of Semantic Enhancement SDS. Our method effectively mitigates semantic coupling issues in SDS, such as generating the tortoise’s tail and head or the deer’s ear and mouth, by providing more focused semantic guidance. In contrast, the semantic suppression method used in CSD and the other semantic enhancement approach proposed by Attend-and-Excite both lead to an unstable optimization.
@@
User Study. We further conducted a user study to evaluate the effectiveness and expressiveness of our method. A Google Form was utilized to assess 1) geometry quality and 2) consistency with text. We recruited 32 participants, of whom 14 are graduate students majoring in media arts, and 18 are company employees specializing in AI content generation. The participants were instructed to choose the preferred renderings of VDM from different methods in randomized order, as shown in Table 2. The results show participants preferred our method by a significant margin.
@@ 4.3 Ablation Study
Effect of Semantic Enhancement SDS. Figure 7 compares the results generated by the original SDS [37], our Semantic Enhancement SDS with Compel, SDS with another semantic enhancement method in Attend-and-Excite (AaE) [7], and CSD [59] with two different annealed weights for the negative prompts: “tortoise tail, tortoise head” and “deer’s ear, deer’s mouth.” We also qualitatively compare these methods by visualizing their performance on semantic decoupling using the same Stable Diffusion for 2D image generation (Figure 8). Compared to semantic suppression using negative prompts and semantic enhancement by AaE, our design of introducing a semantically focused text embedding via Compel to SDS is most effective in resolving the issue of semantic coupling.
Key Insight. As discussed in Section 3.3, SDS can result in semantic coupling when generating sub-object structures, leading to artifacts like the tortoise’s tail and head or the deer’s ear and deer’s mouth. We observed that the semantics represented by negative prompts are also coupled. Therefore, meaningless semantics can emerge when applying semantic suppression in SDS, which leads to unstable optimization. Increasing the weight of negative prompts further reduces the overall quality of generated results. AaE enhances the cross-attention map of specific tokens by continuously updating the latent code. However, AaE is not suitable for the SDS framework because the latent code is affected by different camera poses and random noise at each iteration. This temporal randomness undermines the continuity of the latent code, resulting in unstable optimization and subpar results. In contrast, our method uses Compel to enhance semantics and achieves more effective semantic decoupling. Moreover, the semantically focused text embedding produced by Compel is independent of temporal variation. These properties help mitigate semantic coupling during SDS optimization, leading to high-quality sub-object geometric structures. Meanwhile, our method can also enhance the semantics in the text prompt for surface detail generation (see Appendix A.2).
Figure 8: Visualization of semantic decoupling performance. The images generated by the T2I model qualitatively visualize sub-object semantics. (a) Coupled semantics in positive and negative prompts lead to meaningless distributions when subtracted. (b) Enhanced semantics by Compel achieves stable decoupling. (c) AaE can only ensure that the enhanced semantics are preserved, but cannot decouple negative ones effectively.
Figure 9: Effect of region mask. Region masks can effectively control the pattern of surface details based on text input.
Effect of Region Mask. Figure 9 demonstrates two region masks and their control over surface detail generation given a text prompt. Without a region mask, the results can still match the text prompt but lack a specific pattern. By using a region mask and maintaining an activation ratio of 1/2 throughout the total iterations as a warm-up stage, we achieve a reasonable tradeoff between mask-result alignment and generation diversity (see Appendix A.3).
Effect of VDM Initialization. Our method demonstrates that user-specified VDMs can effectively control the volume and direction of generated geometric structures. Figure 10 shows that the results are high-quality and match the text descriptions well, such as the elf ear and pauldron. As this initializes the Laplacian term and steers the gradient flow in geometric structure generation, users can easily specify an initial VDM or choose a VDM template provided in our framework to generate expressive results.
Figure 10: Effect of VDM initialization. User-specified VDMs can help achieve the intended final effect of geometric structures by initializing the brush’s volume and direction.
@@ 4.4 Applications
Once various VDMs are generated, users can use these brushes to meet diverse creative needs in modeling software. For example, they can apply VDM brushes for mesh stylization and a real-time iterative modeling process.
Local-to-Global Mesh Stylization. Although mesh stylization is a complex task even for professional artists, combining different surface details allows users to achieve stylization quickly (see Appendix C.3).
Coarse-to-Fine Interactive Modeling. Unlike previous methods [4, 60] that require a lengthy optimization process for each edit and result in non-reusable outcomes, our generated VDM brushes can be directly used in modeling software. This enables users to apply the generated brushes easily and interactively (Figure 11).
@@ 5 Conclusion
We have presented Text2VDM, a novel framework for VDM brush generation from text. A VDM is a non-natural 2D image where each pixel stores a 3D displacement vector, making it challenging for existing T2I models to generate. Therefore, we treat VDM generation as diffusion-guided mesh deformation formulated as a form of Sobolev preconditioned optimization. To mitigate semantic coupling issues in SDS, we propose using weighted blending for prompt tokens, achieving high-quality brush generation. Moreover, we introduce two control methods, i.e., region and shape control, to meet customized requirements. The generated VDMs are directly compatible with mainstream modeling software, enabling various applications such as mesh stylization and interactive modeling.
Limitations and Future Work. While our framework can generate high-quality VDM brushes, they may encounter multiview inconsistencies, a common issue introduced by SDS. View-consistent diffusion models like MV2MV [6] may be helpful to further handle this. Additionally, errors of the T2I models may guide wrong 3D generation, such as the wrong number of fingers in the generated skeleton hand. VDMs demonstrate that complex 3D models can be efficiently created using diverse reusable sculpting brushes. Future exploration of 3D generation through the assembly of modular components with similar design principles holds promising research value.
Acknowledgments
This work was partially supported by the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things #2023B1212010007.
Supplementary Material
In this supplementary material, we provide additional details and results that are not included in the main paper due to the space limit.
Appendix A Implementation details
Shape Control. Our framework provides a simple and intuitive interactive interface (Figure 12) to assist users in drawing user-specified VDMs to initialize the base mesh.
Brush Optimization. In each iteration, We randomly sample camera poses in multi-view and render $N$ view normal images at a resolution of 512×512 pixels. We sample the elevation angle as $\phi_{\text{elev}}\sim\mathcal{U}\left(0,\frac{\pi}{3}\right)$, and the azimuth angle as $\phi_{\text{azim}}\sim\mathcal{U}(0,2\pi)$. We set $N=4$ in our experiment. We then feed them into the Stable Diffusion 2.1 to calculate the score distillation sampling (SDS) loss. The generation process runs on a single NVIDIA RTX 4090 GPU with 10, 000 iterations per brush, which takes around 40 minutes.
VDM Baking. For the VDM baking, we write the coordinate values of each vertex of the final mesh into the corresponding pixels, resulting in a three-channel image with a resolution of 512×512 in EXR file format. This EXR image serves as the final VDM output of the generated brush.
A.1 Interface Design for Shape Control
Figure 12: Interface design.
A.2 Details of Weighted Blending
To obtain semantically focused text embeddings for a more precise target distribution to mitigate the semantic coupling issue in SDS, we propose enhancing the semantics of part-related words by applying weighted blending to the tokens in the prompt. Specifically, we assign each word in the prompt a weight $s$ and compute the weighted embedding $e_{w}$ for each word by blending the original text embedding $e$ with the empty text embedding $e_{\phi}$ as follows: $e_{w}=e_{\phi}+s\cdot(e-e_{\phi})$. By concatenating the weighted embeddings of each word in sequence, we obtain the final semantically focused text embedding. For instance, consider the prompt: “A horn++ of a deer”, where ++ indicates that the word “horn” is weighted with a weight of $1.1^{2}$. In the example prompt shown in the main paper, words with yellow underlines represent those weighted with a weight value of $1.21$ during the computation of text embeddings. Notably, the weights used for text embedding computation are separate from the CFG guidance scale applied during the computation of the SDS loss. Our experiment uses a CFG guidance scale of 100. Additionally, our method not only effectively alleviates semantic coupling during geometric structures generation but also enhances specific words in the prompt to produce surface details that better align with the text. For example, Figure 13 demonstrates how enhancing the word ”cannonball” results in a better results.
Figure 13: Effect of Semantic Enhancement SDS on surface details generation.
A.3 Region Control
In our framework, we provide a region mask to restrict mesh deformation to the user-defined region during optimization. By adjusting the activation ratio of the region mask, the final brush effect can effectively match the user’s guidance. When generating surface detail brushes, we activated the region mask during the first half of the total iterations as a warm-up stage, with the specific effects shown in Figure 14. For geometric structures generation, we activated the region mask throughout the entire optimization process to maintain zero values in unused VDM areas.
Figure 14: Region control.
Appendix B Details of Experiment Setting
B.1 Details of Re-framed Paint-it
Paint-it originally uses SDS to optimize a UNet for generating PBR textures. We reframed it to suit our VDM brush generation task. Specifically, in the original Paint-it architecture, a UNet is used to predict a nine-channel output from a fixed 512×512 Gaussian noise image, where every three channels correspond to diffuse, specular, and normal maps, respectively. We modified the UNet to output only three channels, representing a VDM. Each pixel’s three values in the VDM are directly applied to the corresponding mesh vertices for three-axis displacement, resulting in the deformed mesh. We then render the normal map from the deformed mesh to compute the SDS loss. To ensure the quality of the mesh, we introduced regularization terms to facilitate mesh smoothness, such as Laplacian, edge, and normal consistency regularization.
B.2 Base Mesh Initialization
In our experiments for generating geometric structures, to ensure a fair comparison with Text2Mesh and TextDeformer, we used the same initial base mesh and masks as shown in Figure 15. The specific correspondence between prompts and base meshes is detailed in Table 3. Since Paint-it directly generates VDMs, we only used a planar base mesh and do not use masks.
Figure 15: Base mesh initialization.
B.3 Text Prompt of Quantitative Comparison
As described in Section 4.2 of the main paper, we used 40 text prompts for VDM generation and quantitative evaluation, with 20 focusing on surface details and the remaining 20 on geometric structures. The specific text prompts are detailed in Table 3.
Table 3: Text Prompts for VDM Generation.
VDM Type Text Prompt
Surface Details A blanket surface with many flowers
A brick wall surface with neat brick
A broken brick wall surface with many cracks
A broken glass surface like spider web
A broken wall surface damaged by a cannonball
A broken wall surface with many cracks
A cloth surface with a rose pattern
A cloth surface with a sunflower pattern
A torn cloth surface
An arid land surface
A stone surface with many cracks
A broken stone surface with many cracks
A dragon skin surface with many scales
A lion fur surface
A new wooden floor
An aged wooden surface
A rusty metal surface with many cracks
A surface of rusty metal
A skin surface with a scar mended by needle and thread
A skin surface with terrible wounds
Geometric Structures A lip of human (a)
A human spine (a)
A mouth of a monster (a)
A skeleton hand (a)
A tortoiseshell (a)
A round snail shell (a)
An eye of a monster (a)
A dragon wing (b)
An angel wing (b)
An ear of the devil (b)
A dorsal fin of a fish (c)
A horn of a dragon (d)
A goat horn (d)
A horn of the devil (d)
A horn of a deer (d)
An octopus tentacle (d)
An ox horn (d)
A royal pauldron (e)
A beard of a man (e)
A human ear (f)
Figure 16: More application examples of interactive modeling.
Appendix C Additional Results
C.1 More Examples of Semantic Coupling
In this subsection, we present more examples of semantic coupling. As shown in Figure 17, directly using SDS for generating a deer horn leads to the generation of the entire deer head, or generating a beard also results in the generation of the nose. Our method significantly alleviates this semantic coupling issue, enabling the generation of high-quality sub-object structures.
Figure 17: More examples of semantic coupling.
C.2 More Comparison Results
We provide more qualitative results for surface details and geometric structures. Our method can generate higher-quality and more vivid results, as shown in Figure 20 and Figure 21.
Figure 18: More application examples of mesh stylization on a helmet.
Figure 19: More application examples of mesh stylization on a pillar.
C.3 More Application Examples
We also provide more examples of sculpted models using the generated VDM brushes. We enable users to interactively use a variety of generated brushes to sculpt diverse and expressive models from a plain shape.
Figure 20: More comparison results of brushes for surface details.
Figure 21: More comparison results of brushes for geometric structures.